LHC 2013 - latest news

Week 6: Machine Coordinators: G. Arduini, E.B. Holzer

Goal of the week: proton-lead physics and pp run at intermediate energy

p-Pb commissioning plan (John Jowett)

link to old news

Date

Summary

Link to report from morning meeting

Mon 11th February
  • loss maps and asynchronous dump test at injection
08:30 meeting
Sun 10th February
  • 6:04 OP dump of last Pb-p fill
  • preparing for pp run
  • 7:00 - 9:30 access: BPMS IP6, BI in IP4, ADT, ATLAS, CMS
  • interlock settings of BLM, BPMS and RF reverted to high intensity pp settings
  • ADT and feedback settings reverted to pp (bandwidth, gain, gating)
  • ALICE polarity to negative
  • 10:00 - 16:00 2 cycles with probes (tune, chroma, orbit)
  • 17:00 - 19:00 testing interlock BPMS with orbit bumps
  • 19:30 - 20:40 loss maps at injection
  • 20:40 - 4:30 cycle with 5 bunches for: collision set-up, collimator set-up (TCTs and TCLs), vertical RP alignment and data taking, betatron loss maps in collision and asynchronous dump test (RP at 22 sigma)
09:00 meeting
Sat 9th February
  • 02:42 STABLE BEAMS #3541
  • 07:45 Beam dump BPM LSS6
  • 10:40 Stable BEAMS #3542
  • Van der Meer scan in LHCb
  • 15:00-15:20 Adjust. End of fill #3542. Test for measurment of 50 Hz noise at top energy
  • 18:31 STABLE BEAMS 3543
  • ~23:00 Exceeded 30 nb-1 delivered in ALICE/ATLAS/CMS. More than 3nb-1 in 24 h
09:00 meeting
Fri 8th Febraury
  • Fill 338 on 338 with ALICE head on and ATLAS/CMS and LHCb at ~1% of peak luminosity Synchronization problem with SPS delaying fill by ~0.5 h.
  • 02:57 STABLE BEAMS #3538
  • 12:58 BPM LSS6 trigger - Wnd of fill 3538
  • 14:00 Filing for VdM scan ATLAS/LHCb
  • 16:45 STABLE BEAMS #3540
  • 18:30 End of VdM scan in ATLAS. LHCb starts
  • 18:42 RQX.L8 trip. End of fill #3540. Instrumentation problem on DFBX.
  • 22:27 End of access for RQX.L8
  • 23:47 IT.L8 cryo OK
08:30 meeting
Thu 7th Febraury
  • 2:25 fill 3534 dumped by BPMS IP6
  • 3:19 re-injecting
  • 5:24 stable beams fill #3535
  • 10:00-10:30 End of fill test: bump scan at the BSRT position for BSRT calibration. Dump from SIS during the scan
  • 10:31 Dump from SIS during the scan. End of fill #3535
  • 11:30-13:00: Scraping test for energy calibration error evaluation
  • 14:00 filling
  • 14:59 Beam dump during cogging: too large momentum offset excursion for B2 (p) during the process > 10 BPMs out of tolerance
  • 18:17 STABLE BEAMS #3537
  • 23:34 BPM Interlock - End of fill #3537
08:30 meeting
Wed 6th Febraury
  • Disable the batch by batch longitudinal blow-up at injection. Longer bunches from the SPS?
  • 02:36: Large OFB correction at flat-top. Bad BPM was misteriously enabled!!. Beam dump
  • Again issues with the quality of the ion beam in the injection.
  • 05:20 Ramp. This time with acceptable losses (35 %)
  • 06:08 STABLE BEAMS #3533
  • 10:04 LSS6 BPM trigger end of fill #3533
  • 11:00-12:30 Injection checks and energy matching
  • 13:00 lost cryo in S23 --> Access
  • 18:00 Cryo OK
  • 21:22 STABLE BEAMS #3534
08:30 meeting
Tue 5th Febraury
  • 00:04 beam dump after cogging by losses 11R7: orbit drift
  • 3:12 stable beams fill 3527 (338x338)
  • 7:35 dump by BPMS in IP6
  • 10:04 Beam dump when switching ON the OFB. Disabled BPM leading to structure around point 8
  • Filling slowed down by QPS communication problems and poor ion beam quality
  • Increased thresholds for BLMs close to 9 and 11.R7.B1 by 50%.
  • 13:48 STABLE BEAMS 3529
  • 15:57 Trip of Sector 12. (RB.A12 - likely SEU)
  • No switch opening of RCD.A12B2: block this circuit until further investigations
  • LHCb polarity positive
  • 19:39 start filling. Adjustment of injection frequency to minimize capture losses
  • 21:35 PS RF problem during filling
  • 22:30 Beam back. Dump and refill
  • 23:59 Important capture losses at the beginning of the ramp. Beam dump (BLMs in point 3)
08:30 meeting

Mon 4th February
  • 01:00 beam dump" BPM6
  • 02:00 new injection, 192x216 bunches
  • 03:50 stable beams #3523
  • 08:15 End of fill #3523
  • 11:14 Stable Beams #3524 with 192x216 bunches and higher orbit feedback bandwidth (losses at ~20-25% of dump thresoholds)
  • 13:20-13:50 test with local bump in dispersion suppressor R7.B1 to try and spread losses.
  • 15:00 End of fill #3524. BPM6
  • 18:52 Stable beams #3525 with 338x338 bunches. Losses at ~50% of dump threshold at beta* of 1 m.
  • 21:02 beam dump: trip of RQTL9.R3B2
08:30 meeting
Energy calibration